# Technical White Paper (May 2025)

### v1

@@ 1 Introduction
Token consumption and computation costs are predicted to increase exponentially as LLM-based agent applications continue to take over the AI market. Adoption of agentic AI is predicted to increase drastically by the end of this year, with hyper-personalization being the next target on the way to SGI. While hardware strategies are being investigated to mitigate and capture carbon emissions from GPU usage, a new task in classical NLP is emerging [^1] https://tinyurl.com/microsoft-tech-blog  [^2] https://tinyurl.com/ibm-developer-blog: token optimization. In this white paper, we present a new (patent-pending) text representation technique that draws concepts from the fields of data compression, information theory, game theory, and NLP to arrive at a first-of-its-kind word-level semantic text compression algorithm. This algorithm is available for public and commercial use through the Hypernym API, under the product name Mercury333hypernym.ai/pricing.
@@ 3 Methods
  • 1. Parsing and Dart Structuring: The input text is analyzed linguistically and semantically to identify its general concepts and specific details. The content is then represented in a structured format dubbed a dart, which serves as an intermediate representation for compression;
  • 2. Detail Importance Evaluation (Game-Theoretic Analysis): Each detail in the dart is evaluated for its contribution to the overall meaning using Shapley value calculations derived from cooperative game theory. This yields an importance score for every detail.
  • 3. Detail Swapping and Compression Optimization: Guided by the importance scores, the system iteratively removes or replaces less critical details to achieve compression, while attempting to keep the meaning unchanged. This may involve swapping a detail with a more general descriptor (hypernym) plus a minimal indicator of the original detail, or other transformations.
  • 4. Multi-Model Verification: The compressed output is then fed through multiple independent models or analysis modules to verify that it maintains semantic fidelity. If any model detects a significant deviation in meaning or key information, the system can adjust (e.g., reinstate a dropped detail or try an alternative compression for that segment).
  • 5. Output Generation (Dart to Text): Finally, the verified dart structure is converted back into a human-readable compressed text format. The end result is a summary that is highly concise but preserves the essential field of information from the original – hence the name “Field Constrictor,” as it constricts the field of data without strangling its meaning.
@@ 3.1 Dart Structure
  • an attached set of details that provide specific information (that would normally make the text longer if included inline).
@@ 4 Experimental Results
In figure 1, we illustrate the top 15 paragraph token efficiency (upper left corner), the statistical significance of the token efficiency measurements across the entire book (upper right corner), the distribution of compression ratios (lower left corner), and ROUGE-L score comparison between standard and Hypernym-RAG (lower right corner). For a more detailed analysis of these results, see our dedicated benchmark report page [^5] http://max.hypernym.ai/.
@@
In the following figures, we note a trend regardless of model chosen for constriction and reconstruction: the third and fourth passages are consistently harder to compress while retaining the same level of semantic similarity with the original. Note that, here, "passage" and "paragraph" are used interchangeably, for brevity: we applied the dart construction part of the field constriction algorithm 60 times for each paragraph of each passage (see the Appendix). For a custom paragraph test, visit our MCP compressor page [^6] https://v0-react-frontend-with-hypernym.vercel.app/.

### v2

@@ 1 Introduction
Token consumption and computation costs are predicted to increase exponentially as LLM-based agent applications continue to take over the AI market. Adoption of agentic AI is predicted to increase drastically by the end of this year, with hyper-personalization being the next target on the way to SGI. While hardware strategies are being investigated to mitigate and capture carbon emissions from GPU usage, a new task in classical NLP is emerging [^1] https://tinyurl.com/microsoft-tech-blog  [^2] https://tinyurl.com/ibm-developer-blog: token optimization. In this white paper, we present a new (patent-pending) text representation technique that draws concepts from the fields of data compression, information theory, game theory, and NLP to arrive at a first-of-its-kind word-level semantic text compression algorithm. This algorithm is available for public and commercial use through the Hypernym API, under the product name Mercury [^3] https://www.hypernym.ai/pricing.
@@ 3 Methods
  1. 1.
    Parsing and Dart Structuring: The input text is analyzed linguistically and semantically to identify its general concepts and specific details. The content is then represented in a structured format dubbed a dart, which serves as an intermediate representation for compression;
  2. 2.
    Detail Importance Evaluation (Game-Theoretic Analysis): Each detail in the dart is evaluated for its contribution to the overall meaning using Shapley value calculations derived from cooperative game theory. This yields an importance score for every detail.
  3. 3.
    Detail Swapping and Compression Optimization: Guided by importance scores, the system iteratively removes or replaces less critical details to achieve compression while attempting to keep the meaning unchanged. This may involve swapping a detail with a more general descriptor (hypernym) plus a minimal indicator of the original detail, or other transformations.
  4. 4.
    Multi-model Verification: The compressed output is then fed through multiple independent models or analysis modules to verify that it maintains semantic fidelity. If any model detects a significant deviation in meaning or key information, the system can adjust (e.g., reinstate a dropped detail or try an alternative compression for that segment).
  5. 5.
    Output generation (Dart to Text): Finally, the verified dart structure is converted back into a human-readable compressed text format. The end result is a summary that is highly concise, but preserves the essential field of information from the original – hence the name Field Constrictor - as it restricts the field of data without strangling its meaning.
@@ 3.1 Dart Structure
  • a set of details that provide specific information (that would normally make the text longer if included inline).
@@ 4 Experimental Results
In figure 1, we illustrate the token efficiency of the top 15 paragraphs (upper left corner), the statistical significance of the token efficiency measurements throughout the book (upper right corner), the distribution of compression ratios (lower left corner) and the comparison of the ROUGE-L score between standard and Hypernym-RAG (lower right corner). For a more detailed analysis of these results, see our dedicated benchmark report page [^5] http://max.hypernym.ai/.
@@
In the following figures, we note a trend regardless of model chosen for constriction and reconstruction: the third and fourth passages are consistently harder to compress while retaining the same level of semantic similarity with the original. Note that, here, "passage" and "paragraph" are used interchangeably, for brevity: we applied the dart construction part of the field constriction algorithm 60 times for each paragraph of each passage (see the Appendix). For a custom paragraph test, visit our MCP compressor page [^6] https://tinyurl.com/hypernym-mercury.